OpenStudy (anonymous):

Use a graphing device to approximate the negative solution to e^((1/2)(x))=x+2 a. -1.728 b. -1.536 c. -1.498 d. -1.621

OpenStudy (anonymous):

And, Same as exercise 9 except find the positive solution. a. 3.357 b. 3.291 c. 3.516 d. 3.463

OpenStudy (da_scienceman):

U need to plot graphs of e^(1/2x) and x+2 separately. They should both intersect at the point e^(1/2x)=x+2 and hurray, that is the answer

OpenStudy (da_scienceman):

Oh might be u need to solve e^(1/2)-x+2 = 0! I am not sure, just saw your follow up question